A harmonic oscillator has a mass of 0.20kg and unknown spring constant,if the period is T=20g find k

1
Expert's answer
2022-03-22T11:00:32-0400

"\omega=\frac{2\pi}T=0.31~\frac{rad}s,"

"k=\omega^2 m=0.02~\frac N{kg}."